OpenStudy (anonymous):

solve the system by graphing x=-y y+x=6 a. (1,1) b. (1,5) c. infinte number of solutions d. no solution

OpenStudy (anonymous):

hi darth try graphing this with y=-x and y=6-x if the 2 lines intersect, there will be a root or solution,,if not there wil be no solution...

OpenStudy (anonymous):

by solving these two equations.... y=y...that is -x=6-x or -x+x=6 or 0=6

OpenStudy (anonymous):

d is the answer here...no solution

OpenStudy (amistre64):

x+y = 0 x+y = 6 no solution since they are parallel lines that never cross